OpenStudy (anonymous):

A person 100 feet from the base of a tower measured an angle of 60 degrees from the base to the top. a. Find the height of the tower. b. what angle is measure if the person is 200 feet from the base?

OpenStudy (anonymous):

as the height of tower has doubled..the angle will be half of 60 i.e. 30

OpenStudy (dumbcow):

the person makes a right triangle with the base and top of tower height is the opposite side from the angle side 100 ft is adjacent side to angle tan(60) = opposite/adjacent = h/100 h = 100*tan(60) b) tan(x) = h/200 x = inverseTan(h/200), use h found above